how would you solve xy=3
Answer by checkley77(12844)

xy=3
To solve this equation make a table of x * y.
x*y=3
-------
1*3=3
3*1=3
1.5*2=3
2*1.5=3
.3*10=3
10*.3=3
.5*6=3
6*.5=3
1.2*2.5=3
2.5*1.2=3